OpenStudy (anonymous):

Can you walk me through another one? lim e^3x-1/4x x->0

OpenStudy (anonymous):

Do you mean $$\frac 1{4x}$$ ?

OpenStudy (anonymous):



OpenStudy (anonymous):

lim x->0

OpenStudy (anonymous):

Cool do you know L'Hôpital's ( http://en.wikipedia.org/wiki/L'H%C3%B4pital's_rule) rule ?

OpenStudy (anonymous):

some what

OpenStudy (anonymous):

ah yes the \[\infty\]

OpenStudy (anonymous):

because of the limit x to 0 wouldn't the answer be 0?

OpenStudy (anonymous):

$$ \text{ This is } \frac 00 \text{ form a perfect candidate for L'Hôpital's :) } $$

OpenStudy (anonymous):

so the answer is zero? correct?

OpenStudy (anonymous):

No, Please note: The function is is not defined as 0, hence we are taking the limit.

OpenStudy (anonymous):

ugh

OpenStudy (anonymous):

Don't get annoyed its really simple :)

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

I took the derivative of the top and bottom and sub the 0... 3e^3(0) / 4 =3/4?